### v2

# Supplemental Material to ”The Emergence of Vertical Diversity under Disturbance”

Supplemental Material to ”The Emergence of Vertical Diversity under Disturbance”
Yuya Shishikura    Hiroki Ohta
Department of Human Sciences, Obihiro University of Agriculture and Veterinary Medicine, Hokkaido, 080-8555, Japan
(January 12, 2024)
I The equations to determine the stationary solutions of the population dynamics
As mentioned in the main text, based on the model we study, we consider ${\bm{Q}}:=(Q_{1},Q_{2},Q_{3})\in[0,1]^{3}$ where $Q_{h}$ is the density of each state $h$. Then, we may derive the following equation of the population dynamics of ${\bm{Q}}$ for $h=1,3$:
$\frac{dQ_{h}}{dt}=V_{h}({\bm{Q}}),$ (1)
$V_{h}({\bm{Q}})=-\delta_{h,3}rQ_{3}+Q_{2}U\Bigl{(}2\to h\mid{\bm{Q}}\Bigr{)}+\delta_{h,1}r\sum_{h^{\prime}=2}^{3}Q_{h^{\prime}}-Q_{h}U\Bigl{(}h\to 2\mid{\bm{Q}}\Bigr{)}.$
We define $Q_{h}^{*}$ such that $V_{h}({\bm{Q}}^{*})=0$ holds for $h=1,3$.
Let us remind that because of $Q_{1}^{*}+Q_{2}^{*}+Q_{3}^{*}=1$, we have
$Q_{2}^{*}=1-Q_{1}^{*}-Q_{3}^{*},$ (2)
$\sum_{h=1}^{3}hQ_{h}^{*}=-Q_{1}^{*}+Q_{3}^{*}+2.$ (3)
We substitute (2) and (3) into (1) and then obtain the following for $h=1,3$:
$-\delta_{h,3}rQ_{3}^{*}+(1-Q_{1}^{*}-Q_{3}^{*})\frac{(1-r)}{4}\Big{(}1+G_{\beta}^{(h)}(Q_{3}^{*}-Q_{1}^{*})\Big{)}$ (4)
$+\delta_{h,1}r(1-Q_{1}^{*})-Q_{h}^{*}\frac{(1-r)}{4}\Big{(}1-G_{\beta}^{(h)}(Q_{3}^{*}-Q_{1}^{*})\Big{)}=0,$
where
$G_{\beta}^{(h)}(a)=\tanh\beta\Big{(}(\delta_{h,1}-\delta_{h,3})2a+1\Big{)}.$ (5)
I.1 The explicit expressions of $Q_{h}^{*}$ and $\chi_{h}$ at $r\to 0$
Based on (4) for $T>0$, we find the stationary solution of the densities $Q_{1}^{*}$ and $Q_{3}^{*}$ satisfying the following:
$\frac{1}{4}(1-2Q_{1}^{*})(1+\tanh\beta)-\frac{1}{4}Q_{1}^{*}(1-\tanh\beta)=0.$ (6)
Then, taking into account (2), we obtain
$Q_{1}^{*}(T,0)=Q_{3}^{*}(T,0)=\frac{1}{2+\exp(-2\beta)},$ (7)
$Q_{2}^{*}(T,0)=\frac{\exp(-2\beta)}{2+\exp(-2\beta)}.$
Next, we focus on the case of $T=0$ where the following holds:
$\lim_{T\to 0}G_{\beta}^{(h)}(a)=\begin{cases}1&{\rm if}\ \beta\bigg{(}(\delta_{h,1}-\delta_{h,3})2a+1\bigg{)}>0\\ -1&{\rm if}\ \beta\bigg{(}(\delta_{h,1}-\delta_{h,3})2a+1\bigg{)}<0\\ 0&{\rm if}\ \beta\bigg{(}(\delta_{h,1}-\delta_{h,3})2a+1\bigg{)}=0.\end{cases}$ (8)
There are $6$ cases in terms of the possibilities of taking values of $G_{\beta}^{(1)}(a)$ and $G_{\beta}^{(3)}(a)$, each of which can be $-1$, 00, or $1$. Note that if one of $G_{\beta}^{(1)}(a)$ and $G_{\beta}^{(3)}(a)$ is 00, the value of the other is uniquely determined. Indeed, when the two conditions of
$\begin{cases}-2Q_{1}^{*}+2Q_{3}^{*}+1>0,\\ 2Q_{1}^{*}-2Q_{3}^{*}+1>0,\end{cases}$ (9)
hold, based on (4), we obtain
$Q_{3}^{*}=-Q_{1}^{*}+1,$ (10)
where we have used
$\lim_{T\to 0}G_{\beta}^{(h)}(Q_{3}^{*}-Q_{1}^{*})=1,$ (11)
for $h=1,3$. Additionally, the two conditions of (9) lead to
$\frac{1}{4}<Q_{1}^{*}(0,0)<\frac{3}{4},$ (12)
$Q_{3}^{*}(0,0)=1-Q_{1}^{*}(0,0),$
$Q_{2}^{*}(0,0)=0.$
Note that the other $5$ cases for the possibilities of taking values of $G_{\beta}^{(h)}(a)$ give no solutions of equation (4).
In order to compute $\chi_{h}(T,r):=\frac{\partial Q_{h}^{*}}{\partial r}$ for $T>0$, let us perform Taylor series expansion of $Q_{h}^{*}(T,r)$ at $r=0$ for $h=1,3$ as follows:
$Q_{h}^{*}(T,r)=Q_{h}^{*}(T,0)+\chi_{h}(T,0)r+O(r^{2}).$ (13)
Then, substituting this form with (7) into (4), for $O(\beta r)\ll 1$ up to $O(r)$ order, we obtain
$-2\chi_{1}-g_{\beta}\chi_{3}+\frac{4g_{\beta}^{\prime}\beta}{2+g_{\beta}}(\chi_{3}-\chi_{1})+\frac{8}{2+g_{\beta}}=0,$ (14)
$-g_{\beta}\chi_{1}-2\chi_{3}+\frac{4g_{\beta}^{\prime}\beta}{2+g_{\beta}}(\chi_{1}-\chi_{3})-\frac{4g_{\beta}}{2+g_{\beta}}=0,$
where
$g_{\beta}=1+\tanh\beta,$ (15)
$g_{\beta}^{\prime}=\frac{1}{\cosh^{2}\beta},$ (16)
with $g_{0}=1\leq g_{\beta}<\lim_{\beta\to\infty}g_{\beta}=2$.
Solving these equations, we finally obtain
$\chi_{1}(T,0)=4\frac{(2+g_{\beta})(4+g_{\beta}^{2})+4g_{\beta}^{\prime}\beta(2-g_{\beta})}{(2+g_{\beta})^{2}((2+g_{\beta})(2-g_{\beta})+8g_{\beta}^{\prime}\beta)},$ (17)
$\chi_{3}(T,0)=4\frac{-4g_{\beta}(2+g_{\beta})+4g_{\beta}^{\prime}\beta(2-g_{\beta})}{(2+g_{\beta})^{2}((2+g_{\beta})(2-g_{\beta})+8g_{\beta}^{\prime}\beta)},$
where $g_{\beta}^{\prime}\beta\simeq\beta\exp(-2\beta)\to 0$ as $\beta\to\infty$ and $g_{\beta}^{\prime}\beta\simeq\beta\to 0$ as $\beta\to 0$. Note that $\chi_{2}=-\chi_{1}-\chi_{3}$ by definition.
I.2 The explicit expressions of $Q_{h}^{*}$ and $\mu_{h}$ at $T\to 0$
Among $6$ possibilities of taking values of $G_{\beta}^{(1)}$ and $G_{\beta}^{(3)}$, let us consider the situation where the following two conditions hold for $T\ll 1$:
$\begin{cases}-2Q_{1}^{*}+2Q_{3}^{*}+1=0,\\ 2Q_{1}^{*}-2Q_{3}^{*}+1>0.\end{cases}$ (18)
In this case, we need to take the limit of $T\to 0$ carefully. First, we can perform Taylor series expansion of $Q_{h}^{*}(T,r)$ at $T=0$ as
$Q_{h}^{*}(T,r)=Z_{h}+\mu_{h}(0,r)T+O(T^{2}),$ (19)
where $Z_{h}:=Q_{h}^{*}(0,r)$ and $\mu_{h}(T,r):=\frac{\partial Q_{h}^{*}}{\partial T}$. Substituting this form into (4), we obtain, up to $O(T^{0})$ order,
$\frac{(1-r)}{4}(1-Z_{1}-Z_{3})\Bigl{\{}1+\tanh 2(-\mu_{1}+\mu_{3})\Bigr{\}}$
$-\frac{(1-r)}{4}Z_{1}\Bigl{\{}1-\tanh 2(-\mu_{1}+\mu_{3})\Bigr{\}}+r(1-Z_{1})=0,$ (20)
$\frac{1}{2}(1-r)(1-Z_{1}-Z_{3})-rZ_{3}=0,$ (21)
where the second equation is obtained on the condition of (18) corresponding to
$\lim_{T\to 0}G_{\beta}^{(3)}(Q_{3}^{*}-Q_{1}^{*})\to 1.$ (22)
Combining (21) with the first condition of (18) gives
$Z_{1}=$ $Q_{1}^{*}(0,r)=-\frac{1}{4}r+\frac{3}{4},$ (23)
$Z_{3}=$ $Q_{3}^{*}(0,r)=-\frac{1}{4}r+\frac{1}{4},$
$Q_{2}^{*}(0,r)=\frac{r}{2}.$
Let us move onto (20) into which we substitute (23), leading to
$\tanh 2(-\mu_{1}+\mu_{3})$ $=\frac{-r^{2}-10r+3}{(r+3)(1-r)}=:\theta(r).$ (24)
Since $-1<\tanh 2(-\mu_{1}+\mu_{3})<1$ should hold, $r$ satisfies $0<r<r_{0}=-3+2\sqrt{3}$ due to $\lim_{r\to 0}\theta(r)=1$ and $\lim_{r\to r_{0}}\theta(r)=-1$. Thus, it turns out that (23) as the stationary solution is valid only for $0<r<r_{0}$. Note that $\lim_{r\to 0}Q_{1}^{*}(0,r)=\sup Q_{1}^{*}(0,0)=\frac{3}{4}$, $\lim_{r\to 0}Q_{3}^{*}(0,r)=\inf Q_{3}^{*}(0,0)=\frac{1}{4}$, and $\lim_{r\to 0}Q_{2}^{*}(0,r)=Q_{2}^{*}(0,0)=0$ hold.
Next, among the other $5$ possibilities of taking values of $G_{\beta}^{(1)}$ and $G_{\beta}^{(3)}$, when the following two conditions
$\begin{cases}-2Q_{1}^{*}+2Q_{3}^{*}+1<0,\\ 2Q_{1}^{*}-2Q_{3}^{*}+1>0,\end{cases}$ (25)
hold, based on (4) we obtain
$-\frac{1}{2}(1-r)Q_{1}^{*}+r(1-Q_{1}^{*})=0,$ (26)
$\frac{1}{2}(1-r)(1-Q_{1}^{*}-Q_{3}^{*})-rQ_{3}^{*}=0,$
where we have used
$\lim_{T\to 0}G_{\beta}^{(1)}(Q_{3}^{*}-Q_{1}^{*})=-1,$ (27)
$\lim_{T\to 0}G_{\beta}^{(3)}(Q_{3}^{*}-Q_{1}^{*})=1.$ (28)
It reads
$Q_{1}^{*}(0,r)=\frac{2r}{1+r},$ (29)
$Q_{3}^{*}(0,r)=\frac{(1-r)^{2}}{(1+r)^{2}},$
$Q_{2}^{*}(0,r)=\frac{2r(1-r)}{(1+r)^{2}}.$
Substituting (29) into the two conditions of (25) gives
$\left\{\begin{array}[]{l}3r^{2}+10r-1>0,\\ r^{2}+6r-3>0,\end{array}\right.$
which lead to $r_{0}<r\leq 1$. Thus, the obtained solution (29) is valid only for $r_{0}<r\leq 1$. Combining two cases of smaller $r$ and larger $r$ than $r_{0}$, the left-sided limit and the right-sided limit of $Q_{h}^{*}$ as $r\to r_{0}$ coincide. Note that the other $4$ cases for the possibilities of taking values of $G_{\beta}^{(h)}(a)$ give no solutions of equation (4).
In order to compute the response $\mu_{h}(0,r)$, we can perform Taylor series expansion as
$Q_{h}^{*}(T,r)=Q_{h}^{*}(0,r)+\mu_{h}(0,r)T+O(T^{2}).$ (30)
Substituting this form into (4) with (22) for $0<r<r_{0}$, we obtain, up to $O(T)$ order,
$\frac{(1-r)}{4}(1-Q_{1}^{*}-Q_{3}^{*})\Bigl{\{}1+\tanh 2(-\mu_{1}+\mu_{3}+O(T))\Bigr{\}}$
$-\frac{(1-r)}{4}Q_{1}^{*}\Bigl{\{}1-\tanh 2(-\mu_{1}+\mu_{3}+O(T))\Bigr{\}}+r(1-Q_{1}^{*})=0,$ (31)
$(1-r)\mu_{1}+(1+r)\mu_{3}=0.$ (32)
Note that the first equation does not give an explicit exact solution up to $O(T)$ order.
Combining (32) with (24), we obtain
$\mu_{1}(0,r)=-\frac{1+r}{4}\text{artanh }\theta(r),$ (33)
$\mu_{3}(0,r)=\frac{1-r}{4}\text{artanh }\theta(r).$
Note that $\mu_{2}=-\mu_{1}-\mu_{3}$ holds by definition, leading to
$\mu_{2}(0,r)=\frac{r}{2}\text{artanh }\theta(r).$ (34)
Thus, it turns out that $|\mu_{h}|\simeq\log(r-r_{0})$ for each $h$.
Let us move onto the case of $r_{0}<r\leq 1$. By using the same Taylor series expansion, instead of (31), we obtain
$-r\mu_{1}-\frac{1}{2}(1-r)\mu_{1}=0.$ (35)
Then, combining this equation with (32) leads to
$\mu_{1}(0,r)=\mu_{3}(0,r)=\mu_{2}(0,r)=0.$ (36)
I.3 The explicit expressions of $Q_{h}^{*}$ and $\nu_{h}$ at $\beta\to 0$
Based on (4), taking into account
$\lim_{\beta\to 0}G_{\beta}^{(h)}(a)=0,$ (37)
we obtain the following:
$\frac{(1-r)}{4}(1-Q_{1}^{*}-Q_{3}^{*})-\frac{(1-r)}{4}Q_{1}^{*}+r(1-Q_{1}^{*})$ $=0,$ (38)
$\frac{(1-r)}{4}(1-Q_{1}^{*}-Q_{3}^{*})-\frac{(1-r)}{4}Q_{3}^{*}-rQ_{3}^{*}$ $=0.$
Solving these equations, we obtain
$Q_{1}^{*}(\infty,r)=\frac{5r^{2}+10r+1}{3r^{2}+10r+3},$ (39)
$Q_{3}^{*}(\infty,r)=\frac{(1-r)^{2}}{3r^{2}+10r+3},$
$Q_{2}^{*}(\infty,r)=\frac{-3r^{2}+2r+1}{3r^{2}+10r+3},$
which are shown in Fig. 1(a).
(a)Refer to caption
(b)Refer to caption
Figure 1: (a) $Q_{h}^{*}$ as a function of $r$ for $T=\infty$. (b) $D_{\alpha}$ as a function of $r$ for $T=\infty$.
In order to compute the response $\nu_{h}(T,r):=\frac{\partial Q_{h}^{*}}{\partial\beta}$, we perform Taylor series expansion of $Q_{h}^{*}(T,r)$ at $T=\infty$ as
$Q_{h}^{*}(T,r)=Q_{h}^{*}(\infty,r)+\nu_{h}(\infty,r)\beta+O(\beta^{2}).$ (40)
Substituting this form into (4), we obtain
$-\frac{1-r}{4}(\nu_{1}+\nu_{3})-\frac{1-r}{4}\nu_{1}-r\nu_{1}$ $=0,$ (41)
$-\frac{1-r}{4}(\nu_{1}+\nu_{3})-\frac{1-r}{4}\nu_{3}-r\nu_{3}$ $=0.$
Solving these equations, we obtain
$\nu_{1}=\nu_{3}=0,$ (42)
leading to $\nu_{2}=0$ by definition.
Note that by taking the derivative of the stationary solutions from (39) with respect to $r$, we also have
$\chi_{1}(\infty,0)=\frac{20}{9},$ (43)
$\chi_{3}(\infty,0)=-\frac{16}{9},$ (44)
$\chi_{2}(\infty,0)=-\frac{-4}{9},$ (45)
which are consistent with the expression of (17) in the limit of $\beta\to 0$.
II Hill numbers $D_{\alpha}(r)$ as a function of $r$
Let us define $M_{\alpha}:=\log D_{\alpha}$ and ${\bm{Q}}^{*}:=(Q_{1}^{*},Q_{2}^{*},Q_{3}^{*})\in[0,1]^{3}$. Then, the sign of $\partial_{r}M_{\alpha}$ is the same as that of $\partial_{r}D_{\alpha}$. Indeed, we get the following:
$\partial_{r}M_{\alpha}=\alpha(1-\alpha)^{-1}(\sum_{h}(Q_{h}^{*})^{\alpha})^{-1}F_{\alpha}({\bm{Q}}^{*}),$ (46)
$F_{\alpha}({\bm{Q}}^{*})=\sum_{h}(Q_{h}^{*})^{\alpha-1}\chi_{h}.$ (47)
Thus, the nontrivial part determining the sign of $\partial_{r}D_{\alpha}$ is $F_{\alpha}$, which has key information to find out the local maximum point $r_{\ell}(T,\alpha)$ of $D_{\alpha}$. Note that $F_{\alpha}=0$ at $\alpha=1$ holds due to $\sum_{h}\chi_{h}=0$.
II.1 The existence of a local maximum of $D_{\alpha}$ at $T\to 0$ for any $0<\alpha<\infty$
First, let us consider the case of $r<r_{0}$ and $T\to 0$ where we use equations (23) by putting $r=r_{0}$. Then, as shown in Fig. 2(a), we obtain that $F_{\alpha}<0$ for $\alpha>1$, $F_{\alpha}>0$ for $10.199\leq\alpha<$ 1, and $F_{\alpha}<0$ for $\alpha\leq 0.198$ as $r\to r_{0}$. It means that for $\alpha\geq 0.199$, $\partial_{r}D_{\alpha}$ as $r\to r_{0}$ is positive, whereas it is negative for $0<\alpha\leq 0.198$.
Second, let us consider the case of $r>r_{0}$ where we use equations (29) by putting $r=r_{0}$. Then, as shown in Fig. 2(b), it turns out that $F_{\alpha}>0$ for $\alpha>1$ and $F_{\alpha}<0$ for $\alpha<1$. It means that $\partial_{r}D_{\alpha}$ as $r\to r_{0}$ is negative for any $\alpha>0$.
Therefore, combining two cases together, we reach the conclusion that there is a local maximum point $r_{\ell}(0,\alpha)>0$ at $T\to 0$ for any $\alpha>0$. The concrete values of $r_{\ell}(0,\alpha)$ are $r_{0}$ for $\alpha\geq 0.199$, and get smaller than $r_{0}$ between $\alpha=0.198$ and $0.199$. Then, $\lim_{\alpha\to 0}r_{\ell}(0,\alpha)$ is numerically estimated as $0.451\cdots$.
(a)Refer to caption
(b)Refer to caption
Figure 2: (a) $\lim_{r\to r_{0}}F_{\alpha}$ as a function of $\alpha$ for $r<r_{0}$. (b) $\lim_{r\to r_{0}}F_{\alpha}$ a function of $\alpha$ for $r>r_{0}$
II.2 The absence of local maxima of $D_{\alpha}$ at $T\to\infty$ for $0<\alpha<\infty$
Suppose the following condition that for a fixed $T$, $\chi_{1}>0$, $\chi_{2}<0$, $\chi_{3}<0$, and $Q_{1}^{*}>Q_{2}^{*},Q_{3}^{*}$ hold for $0<r\leq 1$. On that condition, we can derive that $(Q_{1}^{*})^{\alpha-1}\chi_{1}$ as a term in $F_{\alpha}$ can be larger than $|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|$ for $\alpha>1$, and smaller than it for $10<\alpha<$ 1. This leads to that for $0<r\leq 1$, $F_{\alpha}>0$ holds for $\alpha>1$, and $F_{\alpha}<0$ holds for $10<\alpha<$ 1, corresponding to that there is no local maximum of $D_{\alpha}$ as a function of $r$ in $0<r\leq 1$ for $\alpha>0$.
The derivation of the above conclusion is as follows. First, it is straightforward to see
$|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|$ $=(Q_{2}^{*})^{\alpha-1}\chi_{1}+(Q_{3}^{*})^{\alpha-1}|\chi_{3}|-(Q_{2}^{*})^{\alpha-1}|\chi_{3}|$
$=(Q_{3}^{*})^{\alpha-1}\chi_{1}+(Q_{2}^{*})^{\alpha-1}|\chi_{2}|-(Q_{3}^{*})^{\alpha-1}|\chi_{2}|,$ (48)
where we have used $\sum_{h}\chi_{h}=0$. Hence, we obtain
$\chi_{1}\min(Q_{2}^{*},Q_{3}^{*})^{\alpha-1}\leq|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|\leq\chi_{1}\max(Q_{2}^{*},Q_{3}^{*})^{\alpha-1}.$ (49)
It means that for $\alpha>1$, the following holds:
$\chi_{1}(Q_{1}^{*})^{\alpha-1}>\chi_{1}\max(Q_{2}^{*},Q_{3}^{*})^{\alpha-1},$ (50)
which immediately leads to $F_{\alpha}>0$. On the other hand, for $10<\alpha<$ 1, the following holds:
$\chi_{1}(Q_{1}^{*})^{\alpha-1}<\chi_{1}\min(Q_{2}^{*},Q_{3}^{*})^{\alpha-1},$ (51)
which immediately leads to $F_{\alpha}<0$. Thus, we reach the conclusion mentioned above.
For $T=\infty$, (39) directly means that the conditions of $\chi_{1}>0,\chi_{2}<0,\chi_{3}<0$, and $Q_{1}^{*}>Q_{2}^{*},Q_{3}^{*}$ hold. Thus, it leads to the absence of local maxima of $D_{\alpha}$ in $0<r\leq 1$ for any $\alpha>0$ as shown in Fig. 1(b).
Note that the conditions mentioned above can be broken, in particular, for sufficiently small $T$, for which the above results are not applied. Nevertheless, using a nullcline method to compute the stationary solutions of (4), we have numerically estimated that there is $\alpha_{0}(T)>0$ such that for $\alpha>\alpha_{0}(T)$, there are no local maxima of $D_{\alpha}$ for $T\geq 0.26$, whereas for $T\leq 0.25$, there is a local maximum of $D_{\alpha}$ for any $\alpha>0$.
III Monte Carlo sampling
III.1 Updating rules
Let us explain how to obtain sample trajectories by Monte Carlo simulations and compare them with the solutions of the derived population dynamics (1). As Monte Carlo simulations, we perform the following procedures:
  1. Initially, prepare a configuration $\bm{x}$ where the state at each site is a value taken randomly from $\{1,2,3\}$ with uniform distribution. That is, the probability of taking each state is $1/3$.
    Next, repeat the steps from $1$ to $3$:
  2. 1.
    Choose randomly a name $\ell\ (1\leq\ell\leq N)$ of an agent with uniform distribution. That is, the probability of taking each name is $1/N$.
  3. 2.
    With the probability $r$, the chosen agent $\ell$ is disturbed, making the transition from state $x_{\ell}$ to $1$.
  4. 3.
    If the chosen agent $\ell$ is not disturbed, choose either positive or negative direction with probability $1/2$.
    1. (a)
      If positive direction is chosen and $x_{\ell}\neq 3$, make the transition from state $x_{\ell}$ to $x_{\ell}+1$ with the probability
      $\frac{1}{2}\Bigg{(}1+\tanh{\beta\Big{(}2(x_{\ell}-\frac{1}{N}\sum_{k\neq\ell}x_{k})}+1\Big{)}\Bigg{)}.$ (52)
    2. (b)
      If negative direction is chosen and $x_{\ell}\neq 1$, make the transition from state $x_{\ell}$ to $x_{\ell}-1$ with the probability
      $\frac{1}{2}\Bigg{(}1+\tanh\beta\Big{(}-2(x_{\ell}-\frac{1}{N}\sum_{k\neq\ell}x_{k})+1\Big{)}\Bigg{)}.$ (53)
As shown in Fig. 3, we have confirmed that Monte Carlo simulations have good agreement with the solution of the population dynamics (1).
Figure 3: The Monte Carlo (MC) simulation with $N=1000$ and the solutions of the population dynamics (1). $T=1$, $r=0.2$.
III.2 Finite-size fluctuations
Let us characterize the finite size fluctuations in the density of each state by variance $\sigma_{hh}$, covariance $\sigma_{hh^{\prime}}$, and correlation coefficient $C_{hh^{\prime}}$ for $h,h^{\prime}\in\{1,2,3\}$, which are defined as
$\sigma_{hh^{\prime}}:=\frac{N}{\tau-t_{0}}\sum_{t=t_{0}}^{\tau}\Delta_{h}(t)\Delta_{h^{\prime}}(t),$ (54)
$\Delta_{h}(t):=\hat{Q}_{h}(t)-Q_{h}^{*},$ (55)
$C_{hh^{\prime}}:=\frac{\sigma_{hh^{\prime}}}{\sqrt{\sigma_{hh}}\sqrt{\sigma_{h^{\prime}h^{\prime}}}},$ (56)
where $\hat{Q}_{h}(t)$ is a sample trajectory of density of state $h$ obtained from Monte Carlo simulation at time $t$. The time unit is one Monte Carlo step per site ($N$ steps). We set $t_{0}=100$ because the typical trajectories of $\hat{Q}_{h}(t)$ with $r\geq 10^{-2}$ and $T=10^{-3}$ crosses the stationary solution much before $t=t_{0}=100$. Indeed, for $r\leq 10^{-3}$ and $T=10^{-3}$, the dynamics gets very slow and the typical trajectories do not cross the stationary solutions by $t=t_{0}=100$.
Figure 4: The variance $\sigma_{hh}$ and correlation coefficients $C_{hh^{\prime}}$ averaged over $10$ samples with different trajectories. $T=10^{-3}$, $t_{0}=100$, $\tau=150$, $N=10^{2},10^{3},10^{4}$.
As shown in Fig. 4, we have performed Monte Carlo simulations with different systems size $N$. If we straightforwardly extrapolate the behaviors for larger system size, there seem to be singular behaviors in the three types of variance and the three types of correlation function in the large size limit near $r=r_{0}$ for $r<r_{0}$. In particular, it is notable to point out that three correlation coefficients in the large size limit seem to have perfect correlation, corresponding to $|C_{hh^{\prime}}|=1$, for only $r<r_{0}$.
Thus, the results obtained by Monte Carlo simulations suggest that there is a universal curve presumably independent of $N$. Further, in the larger system sizes, it is more obvious that $C_{hh^{\prime}}$ for $r<r_{0}$ show U-shaped curves. It is notable that the bottoms of the U-shaped curves are close to $r=-5+2\sqrt{7}\simeq 0.291\cdots$, which is the point where $\mu_{h}$ changes its sign. Nevertheless, more works with sufficiently large samples need to be done to obtain more convincing data, which remain to be found as another work in the future.